Please help me solve the following probability question
The average length of stay in a hospital is useful for planning purposes. Suppose that the following is the distribution of the length of stay in a hospital after a minor operation.
x = 1, 2, 3, 4, 5
P(x) = 0.6, 0.1, 0.04, 0.02, c
What is the average length of stay in the hospital?

Answer by jim_thompson5910(35256)

Given Distribution
X12345
P(X)0.60.10.040.02c


First we need to know the value of c. The probabilities must add to 1 for this to be a true probability distribution

0.6+0.1+0.04+0.02+c = 1

0.76+c = 1

0.76+c-0.76 = 1-0.76

c = 0.24

So the probability distribution updates to
X12345
P(X)0.60.10.040.020.24


----------------------------------------------------------------------------------

To find the average, multiply the probabilities with their corresponding x values. Then add up those products.

1*0.6+2*0.1+3*0.04+4*0.02+5*0.24 = 2.2

----------------------------------------------------------------------------------
----------------------------------------------------------------------------------

Final Answer: The average length of stay in the hospital is 2.2 days